a student takes a 20 question multiple choice exam with 5 choices for each question and guesses on each question. find the probability of guessing at least 15 out of the 20 correctly.
Answer by stanbon(75887)

a student takes a 20 question multiple choice exam with 5 choices for each question and guesses on each question. find the probability of guessing at least 15 out of the 20 correctly.
-----------------------
It is a binomial problem with n=20, p=1/5 ; 15<= x <=20
--------------------------
Prob = 1 - binomcdf(20,1/5,14) = 0.0000001803
====================================================
Cheers,
Stan H.